# Electric field induction in quark-gluon plasma due to thermoelectric effects

### v1

@@ 3.1.1 Analytical solution for massless non-interacting system
where, $\Gamma(s)$ is the Gamma function. Now, using the relation $\eta(s)=1-2^{(1-s)}\zeta(s)$, Eqs. (3.1) can be write in terms of the $\zeta(s)$ as,
@@
$L_{23}$ $=\sum_{i}q_{i}~{}\frac{g_{i}\tau_{R}^{i}T^{2}}{6\pi^{2}}\Big{(}\frac{b_{i}^{3}\mu_{B}^{3}}{T^{2}}+6b_{i}\mu_{B}\zeta(2)-2b_{i}h\zeta(2)-\frac{b_{i}^{3}\mu_{B}^{2}h}{T^{2}}\Big{)}\left(1-\tau_{R}^{i}\frac{d}{dt}\right).$ (30)
Note that, $L_{23}$ is proportional to electric charge $q_{i}$. Degeneracy is the same for all three flavors, $g_{i}=6$ (for a spin and color degeneracy). $\tau_{R}^{i}=\tau_{R}$ (say) is the same for all the flavors for massless and non-interacting systems of particles. Therefore, the charge of the three flavors (up, down, and strange) adds up to zero, i.e., $L_{23}=0$. As a result, the Seebeck coefficient for the massless non-interacting system is
@@
3.1.2 Numerical solution for QGP EoS
Refer to captionRefer to caption
@@ 3.1.2 Numerical solution for QGP EoS
We use a quasiparticle model formulated by Gorenstein and Yang Gorenstein:1995vm for the numerical estimation. It is a phenomenological model where the lattice QCD equation of state for QGP is achieved by considering the thermal masses of the partons. The thermal mass $m(T)$ arises from the interactions among the partons. The thermodynamic consistency is achieved by introducing a bag constant arising from vacuum energy Srivastava:2010xa . The dispersion relation of the particle having energy $\omega_{i}$ and momentum $k_{i}$ is $\omega_{i}^{2}(k_{i},T)=k_{i}^{2}+m_{i}^{2}(T)$. Where $m_{i}$ is the total effective mass of ith quark flavor and can be parameterized as
@@
$m_{g}=\frac{N_{c}}{6}g^{2}(T)T^{2}\left(1+\frac{N_{f}+\frac{\mu_{B}^{2}}{\pi^{2}T^{2}}}{6}\right).$ (34)
$N_{c}$ represents the number of color degrees of freedom and $g^{2}(T)=4\pi\alpha_{s}(T)$, $\alpha_{s}(T)$ is running coupling constant. For more details of the model, one can follow references Gorenstein:1995vm ; Srivastava:2010xa ; Bannur:2006hp ; Peshier:1995ty ; Peshier:1999ww . Note that, for the estimation of thermoelectric coefficients and the induced electric field, gluons contribute only to the enthalpy of the system. Now, for the relaxation time of quarks, we use a momentum-independent expression obtained for QCD matter Hosoya:1983xm
@@
In Fig. (2), we plot the Seebeck coefficient ($S$) as a function of proper time ($\tau$) at $\mu_{B}$ = 0.3 GeV in head-on collisions (or $eB=0$) for the static and evolving case of QGP. Plots in the left panels are obtained in Bjorken flow with the ideal hydrodynamic cooling rate, which in the right panels are for Gubser flow. In both cases, a decreasing (magnitude) trend of $S$ with time is observed, and it approaches the massless non-interacting (black dotted line) limit. Note that the negative sign in the value of $S$ implies that the direction of the induced electric field is opposite to the temperature gradient. With time, the temperature of the medium falls off, and $S$ reduces in both the hydrodynamic pictures. $S$ for Gubser flow reduces comparatively faster because of its cooling rate. In both the hydrodynamic pictures, $S$ beyond a certain time where $T<0.150$ GeV is not justified in the used quasiparticle model. From Eq. (27), $S$ is the ratio of two different quantities; for the static case where integral $L_{23}$ reduces to $L_{2i}$, the coefficient $S$ becomes independent of relaxation time $\tau_{R}$. This happens because we consider the same relaxation time for all the quarks regardless of their flavor. Whereas, for the evolving picture, $S$ shows its dependence on $\tau_{R}$ through integral Ł$\L_{23}$. In the bottom panels of Fig. (2), the percentage deviation of static from the evolving picture is plotted against $\tau$ to quantify the cooling effect. In both the hydrodynamic cases, $S$ is enhanced due to temperature evolution. The effect is high in the early times and decreases with time. Moreover, this is higher in the Gubser flow case, which shows the impact is sensitive to the cooling rates.
3.2 Induced electric field in isotropic system
Refer to captionRefer to captionRefer to caption
@@ 3.2 Induced electric field in isotropic system
Figure 3: Electric field induced in the QGP in the head-on collisions.
@@
To estimate the induced electric field ($\vec{E}$) in the head-on collisions, we considered baryon chemical potential $\mu_{B}=0.3$ GeV, attained at RHIC energies. Here, we used the Gubser cooling rate for the temperature gradient. From Gubser flow, $T=T(\tau,~{}r)$, with $r=\sqrt{x^{2}+y^{2}}$. Due to Gubser symmetry considerations as mentioned in subsection (2.2), $T$ is symmetric in the transverse ($x-y$) plane and remains invariant along the longitudinal direction or $z$ axis. Here, we used the Seebeck coefficient calculated from the Bjorken flow for the case of evolving QGP. In head-on collisions, there is no spectator current; the magnetic field is not produced during the collision. In this paper, by the terms “isotropy” and “anisotropy,” we indicate momentum isotropy or anisotropy due to magnetic field only. We will discuss the induced electric field created in an isotropic QGP (without an external magnetic field) due to the thermoelectric effect.
@@
In Fig. (3), we have shown the vector plot of the induced electric field in the transverse plane for head-on collision. In an isotropic medium, the thermoelectric transport coefficient matrix is isotropic as obtained in Eq. (3.1), only diagonal elements, Seebeck coefficient ($S$) are non-zero. Due to symmetry consideration, $T$ and its gradient are symmetric in the transverse ($x-y$) plane. Moreover, because of boost invariant symmetry along $z$ axis, $\frac{\partial T}{\partial z}=0$. Therefore, only $x$ and $y$ components of the electric field survive, and $\vec{E}$ is symmetric in the transverse plane. Time evolution of $\vec{E}$ is shown by three panels in the Fig. (3), from left to right $\tau=0.5,~{}1,~{}3$ fm. The direction and magnitude of $\vec{E}$ are represented in the figures by the arrow tip and its length, respectively. Magnitude is also colored for clarity; the color bar on the top of the figures is given for magnitude in the unit of $m_{\pi}^{2}$. $\vec{E}$ is zero at the center (as $\vec{\nabla}T$ vanishes) and directed radially outward with increasing magnitude. At $\tau=0.5$ fm, the induced field as high as $eE=0.6~{}m_{\pi}^{2}$, and its strength decreases as the medium evolves in time. The direction of $\vec{E}$ can be understood as follows. From Eq. (3.1), $E_{x}=S\frac{dT}{dx}$, $E_{y}=S\frac{dT}{dy}$. From the right panel of Fig. (2), $S$ is negative at the early time. $\vec{\nabla}T$ is also negative, as $T$ decreases as we go away from the center ($x=y=0,~{}{\rm or},~{}r=0$). This explains the direction of $\vec{E}$. Value of $S$, $\vec{\nabla}T$, and hence $\vec{E}$ is sensitive to the cooling rate. As an analytical expression of the cooling rate is required for the calculation of thermoelectric coefficients and induced field, we used Gubser flow. However, in (3+1)D hydrodynamic simulations, one can incorporate the current study to estimate induced fields in QGP produced in heavy-ion collisions.
4 Peripheral collision
@@
In this section, we calculate the thermoelectric coefficients of an evolving relativistic fluid in the presence of an external time-varying magnetic field. For detailed calculations, see appendix A.
4.1 Magneto-Seebeck and Nernst coefficients in evolving QGP
@@
4.1.1 Without Landau quantization
In the presence of a magnetic field, to find the expression of the deviated part of distribution function $\delta f_{i}$, which is driven by the temperature gradient, we use the RTA method with a Lorentz force term in BTE. In the presence of an external electromagnetic field, BTE under RTA can be expressed as PhysRevD.102.016016
@@ 4.1.1 Without Landau quantization
where $\tau^{i}_{R}$ is the relaxation time of the particle. We consider a time-varying electromagnetic field of the form  Satow:2014lia ; Hongo:2013cqa
@@
where $B_{0},~{}E_{0}$ are the magnitudes of the initial fields having decay parameters of $\tau_{B}$ and $\tau_{E}$, respectively, and $\tau$ is the proper time. The deviation of the distribution function from equilibrium is driven by the electromagnetic field, so leading order contribution in $\delta f_{i}$ can be PhysRevD.104.094037
@@
$\vec{\Omega}=$ $~{}\alpha_{1}\vec{E}+\alpha_{2}\dot{\vec{E}}+\alpha_{3}\vec{\nabla}T+\alpha_{4}\vec{\nabla}\dot{T}+\alpha_{5}(\vec{\nabla}T\times\vec{B})+\alpha_{6}(\vec{\nabla}T\times\dot{\vec{B}})$
$+\alpha_{7}(\vec{\nabla}\dot{T}\times\vec{B})+\alpha_{8}(\vec{E}\times{\vec{B}})+\alpha_{9}(\vec{E}\times{\dot{\vec{B}}})+\alpha_{10}(\dot{\vec{E}}\times{\vec{B}}),$ (39)
@@
In the current work, we introduce the cooling rate using hydro-dynamical theories and study the thermoelectric response of evolving QGP in the presence of a slowly varying magnetic field, where we can consider that the inverse of cyclotron frequency is approximately equal to the magnetic field decay parameter, i.e., $\tau_{B}\approx\frac{\omega_{i}}{q_{i}B}$. Therefore, we can solve Eq. (36) using Eqs. (38) and (4.1.1), and get the expression of $\alpha_{j}$’s. After getting the expressions for $\alpha_{j}$’s, the simplified form of $\delta f_{i}$ is(see Appendix A),
$\delta f_{i}$ $=\frac{-q_{i}\tau_{R}^{i}}{(1+\chi_{i})(1+\chi_{i}+\chi_{i}^{2})}\Big{[}\Big{\{}(1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{\}}(\vec{v_{i}}.\vec{E})\pm\Big{\{}\chi_{i}(1+\chi_{i})$
$+\frac{\chi_{i}^{2}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{\}}\Big{(}\vec{v_{i}}.(\vec{E}\times\hat{b})\Big{)}\Big{]}+\frac{(\omega_{i}-{\rm b}_{i}h)\tau_{R}^{i}}{T(1+\chi_{i})(1+\chi_{i}+\chi_{i}^{2})}\Big{[}(1+\chi_{i})\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{T}\Big{)}$
$-\tau_{R}^{i}\frac{(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{\dot{T}}\Big{)}\pm\chi_{i}(1+\chi_{i})\Big{(}\vec{v_{i}}\cdot(\vec{\nabla}{T}\times\hat{b})\Big{)}\mp\tau_{R}^{i}\frac{\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}$
$\Big{(}\vec{v_{i}}\cdot(\vec{\nabla}{\dot{T}}\times\hat{b})\Big{)}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (40)
@@
$j^{l}$ $=\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}|\vec{k}_{i}|}{(2\pi)^{3}}v_{i}^{2}\frac{\tau_{R}^{i}}{(1+\chi_{i}+\chi_{i}^{2})(1+\chi_{i})}\Big{[}-q_{i}\Big{\{}\Big{(}(1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{)}\delta^{lk}E^{k}$
±$\pm\Big{(}\chi_{i}(1+\chi_{i})+\frac{\chi_{i}^{2}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{)}\epsilon^{ljk}h^{j}E^{k}\Big{\}}+\frac{(\omega_{i}-{\rm b}_{i}h)}{T}\Big{\{}(1+\chi_{i})\delta^{lk}\frac{\partial{T}}{\partial x^{k}}$
$-\tau_{R}^{i}\frac{(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\delta^{lk}\frac{\partial{\dot{T}}}{\partial x^{k}}\pm\chi_{i}(1+\chi_{i})\epsilon^{ljk}h^{j}\frac{\partial{T}}{\partial x^{k}}\mp\tau_{R}^{i}\frac{\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}\epsilon^{ljk}h^{j}\frac{\partial{\dot{T}}}{\partial x^{k}}\Big{\}}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (41)
@@
$j^{x}=$ $\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}|\vec{k}_{i}|}{(2\pi)^{3}}v_{i}^{2}\frac{\tau_{R}^{i}q_{i}}{(1+\chi_{i}+\chi_{i}^{2})(1+\chi_{i})}\Big{\{}\Big{(}\frac{(1+\chi_{i}^{2})+\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{)}E^{x}\pm\chi_{i}$
$\Big{(}\frac{(1+\chi_{i}^{2})(1+\chi_{i})+\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{)}(E^{z}-E^{y})\Big{\}}\Big{(}-\frac{\partial f^{0}_{i}}{\partial\omega_{i}}\Big{)}+\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}|\vec{k}_{i}|}{(2\pi)^{3}}v_{i}^{2}$
$\frac{\tau_{R}^{i}(\omega_{i}-{\rm b}_{i}h)}{T(1+\chi_{i}+\chi_{i}^{2})(1+\chi_{i})}\Big{\{}(1+\chi_{i})\frac{\partial T}{\partial x}-\tau_{R}^{i}\frac{(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\frac{\partial\dot{T}}{\partial x}\pm\chi_{i}(1+\chi_{i})\Big{(}\frac{\partial T}{\partial z}-\frac{\partial T}{\partial y}\Big{)}$
@@
we can write Eq. (4.1.1) as
$j^{x}$ $=\sum_{i}q_{i}^{2}L_{1i}E_{x}-\sum_{i}q_{i}^{2}L_{2i}E_{y}+\sum_{i}q_{i}^{2}L_{2i}E_{z}-\frac{1}{T}\sum_{i}q_{i}\Big{(}L_{3i}-L_{4i}\frac{d}{dt}\Big{)}\frac{dT}{dx}+\frac{1}{T}\sum_{i}q_{i}$
$\Big{(}L_{5i}-L_{6i}\frac{d}{dt}\Big{)}\frac{dT}{dy}-\frac{1}{T}\sum_{i}q_{i}\Big{(}L_{5i}-L_{6i}\frac{d}{dt}\Big{)}\frac{dT}{dz}~{}.$ (44)
@@
Where, $\sigma_{el}$ = $\sum_{i}q_{i}^{2}L_{1i}$ and $\sigma_{H}$ = $\sum_{i}q_{i}^{2}L_{2i}$ are the Ohmic and the Hall components of electrical conductivity, respectively. Further, we defined the intergrals $L_{34}$ = $\sum_{i}q_{i}\Big{(}L_{3i}-L_{4i}\frac{d}{dt}\Big{)}$ and $L_{56}$ = $\sum_{i}q_{i}\Big{(}L_{5i}-L_{6i}\frac{d}{dt}\Big{)}$. To solve the Eqs. (45), we rearrange them in the matrix form of
$\mathbb{\bm{\sigma}~{}E=L~{}X},$ (46)

### v2

@@ III.1.1 Analytical solution for massless noninteracting system
where, $\Gamma(s)$ is the gamma function. Now, using the relation $\eta(s)=1-2^{(1-s)}\zeta(s)$, Eqs. (III.1) can be write in terms of the $\zeta(s)$ as,
@@
$L_{23}$ $=\sum_{i}q_{i}~{}\frac{g_{i}\tau_{R}^{i}T^{2}}{6\pi^{2}}\Big{(}\frac{b_{i}^{3}\mu_{B}^{3}}{T^{2}}+6b_{i}\mu_{B}\zeta(2)-2b_{i}h\zeta(2)$
$-\frac{b_{i}^{3}\mu_{B}^{2}h}{T^{2}}\Big{)}\left(1-\tau_{R}^{i}\frac{d}{dt}\right).$ (30)
Note that, $L_{23}$ is proportional to electric charge $q_{i}$. Degeneracy is the same for all three flavors, $g_{i}=6$ (for a spin and color degeneracy). $\tau_{R}^{i}=\tau_{R}$ (say) is the same for all the flavors for massless and noninteracting systems of particles. Therefore, the charge of the three flavors (up, down, and strange) adds up to zero, i.e., $L_{23}=0$. As a result, the Seebeck coefficient for the massless noninteracting system is
@@
III.1.2 Numerical solution for QGP EoS
@@
We use a quasiparticle model formulated by Gorenstein and Yang [59] for the numerical estimation. It is a phenomenological model where the lattice QCD equation of state for QGP is achieved by considering the thermal masses of the partons. The thermal mass $m(T)$ arises from the interactions among the partons. The thermodynamic consistency is achieved by introducing a bag constant arising from vacuum energy [60]. The dispersion relation of the particle having energy $\omega_{i}$ and momentum $k_{i}$ is $\omega_{i}^{2}(k_{i},T)=k_{i}^{2}+m_{i}^{2}(T)$. Where $m_{i}$ is the total effective mass of ith quark flavor and can be parametrized as
@@
$m_{g}^{2}(T,\mu_{B})=\frac{N_{c}}{6}g^{2}(T)T^{2}\left(1+\frac{N_{f}+\frac{\mu_{B}^{2}}{\pi^{2}T^{2}}}{6}\right).$ (34)
$N_{c}$ represents the number of color degrees of freedom and $g^{2}(T)=4\pi\alpha_{s}(T)$, $\alpha_{s}(T)$ is running coupling constant. We have used a One-loop leading-order running coupling constant [61, 62], which is similar to the one used in lattice calculations for QCD. For more details of the model, one can follow references [59, 60, 63, 64, 65, 66]. Note that, for the estimation of thermoelectric coefficients and the induced electric field, gluons contribute only to the enthalpy of the system. Now, for the relaxation time of quarks, we use a momentum-independent expression obtained for QCD matter [67]
@@
In Fig. (2), we plot the Seebeck coefficient ($S$) as a function of proper time ($\tau$) at $\mu_{B}$ = 0.3 GeV in head-on collisions (or $eB_{0}=0$) for the static and evolving picture of QGP. Plots in the left panels are obtained in Bjorken flow with the ideal hydrodynamic cooling rate, which in the right panels are for Gubser flow. In both the flow cases, we observed that the magnitude of coefficient $S$ decreases with an increase in the values of $\tau$. In both plots, we observe that coefficient $S$ approaches zero at later stages of evolution, whereas, for the limiting case of massless partons (black dotted line), the value of $S$ is zero throughout the evolution. Note that the negative sign in the value of $S$ implies that the direction of the induced electric field is opposite to the temperature gradient. With time, the temperature of the medium falls off, and $S$ reduces in both the hydrodynamic pictures. $S$ for Gubser flow reduces comparatively faster because of its cooling rate. From Eq. (27), $S$ is the ratio of two different quantities; for the static picture where integral $L_{23}$ reduces to $L_{2i}$, the coefficient $S$ becomes independent of relaxation time $\tau_{R}$. This happens because we consider the same relaxation time for all the quarks regardless of their flavor. Whereas, for the evolving picture, $S$ shows its dependence on $\tau_{R}$ through integral Ł$\L_{23}$. In the bottom panels of Fig. (2), the percentage deviation of static from the evolving picture is plotted against $\tau$ to quantify the cooling effect. In both the hydrodynamic cases, $S$ is enhanced due to temperature evolution. The effect is high in the early times and decreases with time. Moreover, this is higher in the Gubser flow case, which shows the impact is sensitive to the cooling rates.
III.2 Induced electric field in isotropic system
@@
Figure 3: Time evolution of the induced electric field in the QGP in the head-on collisions with $eB_{0}=0$. Left: at $\tau=0.5$ fm, middle: at $\tau=1$ fm, and right: at $\tau=3$ fm.
@@
To estimate the induced electric field ($\vec{E}$) in the head-on collisions, we considered baryon chemical potential $\mu_{B}=0.3$ GeV, attained at RHIC energies. Here, we used the Gubser cooling rate for the temperature gradient. From Gubser flow, $T=T(\tau,~{}r)$, with $r=\sqrt{x^{2}+y^{2}}$. Due to Gubser symmetry considerations as mentioned in subsection (II.2), $T$ is symmetric in the transverse ($x-y$) plane and remains invariant along the longitudinal direction or $z$ axis. Here, we used the Seebeck coefficient calculated from the Bjorken flow for the case of evolving QGP. In head-on collisions, there is no spectator current; the magnetic field is not produced during the collision. In this paper, by the terms “isotropy” and “anisotropy,” we indicate momentum isotropy or anisotropy due to magnetic field only. We will discuss the induced electric field created in an isotropic QGP (without an external magnetic field) due to the thermoelectric effect.
@@
In Fig. (3), we have shown the vector plot of the induced electric field in the transverse plane for head-on collision. In an isotropic medium, the thermoelectric transport coefficient matrix is isotropic as obtained in Eq. (III.1), only diagonal elements, Seebeck coefficient ($S$) are nonzero. Due to symmetry consideration, $T$ and its gradient are symmetric in the transverse ($x-y$) plane. Moreover, because of boost invariant symmetry along $z$ axis, $\frac{\partial T}{\partial z}=0$. Therefore, only $x$ and $y$ components of the electric field survive, and $\vec{E}$ is symmetric in the transverse plane. Time evolution of $\vec{E}$ is shown by three panels in the Fig. (3), from left to right $\tau=0.5,~{}1,~{}3$ fm. The direction and magnitude of $\vec{E}$ are represented in the figures by the arrow tip and its length, respectively. Magnitude is also colored for clarity; the color bar on the top of the figures is given for magnitude in the unit of $m_{\pi}^{2}$. $\vec{E}$ is zero at the center (as $\vec{\nabla}T$ vanishes) and directed radially outward with increasing magnitude. At $\tau=0.5$ fm, the induced field is as high as $eE\approx 0.8~{}m_{\pi}^{2}$, and its strength decreases as the medium evolves in time. The direction of $\vec{E}$ can be understood as follows. From Eq. (III.1), $E_{x}=S\frac{dT}{dx}$, $E_{y}=S\frac{dT}{dy}$. From the right panel of Fig. (2), $S$ is negative at the early time. $\vec{\nabla}T$ is also negative, as $T$ decreases as we go away from the center ($x=y=0,~{}{\rm or},~{}r=0$). This explains the direction of $\vec{E}$. Value of $S$, $\vec{\nabla}T$, and hence $\vec{E}$ is sensitive to the cooling rate. As an analytical expression of the cooling rate is required for the calculation of thermoelectric coefficients and induced field, we used Gubser flow. However, in (3+1)D hydrodynamic simulations, one can incorporate the current study to estimate induced fields in QGP produced in heavy-ion collisions.
IV Electric field induction in Peripheral collisions
@@
In peripheral heavy-ion collisions, a strong transient magnetic field is produced. In this section, we calculate an evolving relativistic fluid’s thermoelectric coefficients and induced electric field in an external time-varying magnetic field. For detailed calculations, see Appendix A.
IV.1 Magneto-Seebeck and Nernst coefficients in evolving QGP
@@
IV.1.1 Without Landau quantization
In the presence of a magnetic field, to find the expression of the deviated part of distribution function $\delta f_{i}$, which is driven by the temperature gradient, we use the RTA method with a Lorentz force term in BTE. In the presence of an external electromagnetic field, BTE under RTA can be expressed as [15]
@@ IV.1.1 Without Landau quantization
where $\tau^{i}_{R}$ is the relaxation time of the particle. We consider a time-varying electromagnetic field of the form  [51, 55]
@@
where $B_{0},~{}E_{0}$ are the magnitudes of the initial fields having decay parameters of $\tau_{B}$ and $\tau_{E}$, respectively, and $\tau$ is the proper time. The deviation of the distribution function from equilibrium is driven by the electromagnetic field, so leading order contribution in $\delta f_{i}$ can be [58]
@@
$\vec{\Omega}=$ $~{}\alpha_{1}\vec{E}+\alpha_{2}\dot{\vec{E}}+\alpha_{3}\vec{\nabla}T+\alpha_{4}\vec{\nabla}\dot{T}+\alpha_{5}(\vec{\nabla}T\times\vec{B})$
$+\alpha_{6}(\vec{\nabla}T\times\dot{\vec{B}})+\alpha_{7}(\vec{\nabla}\dot{T}\times\vec{B})+\alpha_{8}(\vec{E}\times{\vec{B}})$
$+\alpha_{9}(\vec{E}\times{\dot{\vec{B}}})+\alpha_{10}(\dot{\vec{E}}\times{\vec{B}}),$ (39)
@@
In the current work, we introduce the cooling rate using hydro-dynamical theories and study the thermoelectric response of evolving QGP in the presence of a slowly varying magnetic field, where we can consider that the inverse of cyclotron frequency is approximately equal to the magnetic field decay parameter, i.e., $\tau_{B}\approx\frac{\omega_{i}}{q_{i}B}$. Therefore, we can solve Eq. (36) using Eqs. (38) and (IV.1.1), and get the expression of $\alpha_{j}$’s. After getting the expressions for $\alpha_{j}$’s, the simplified form of $\delta f_{i}$ is(see Appendix A),
$\delta f_{i}$ $=\frac{-q_{i}\tau_{R}^{i}}{(1+\chi_{i})(1+\chi_{i}+\chi_{i}^{2})}\Big{[}\Big{\{}(1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{\}}$
$(\vec{v_{i}}.\vec{E})\pm\Big{\{}\chi_{i}(1+\chi_{i})+\frac{\chi_{i}^{2}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{\}}\Big{(}\vec{v_{i}}.(\vec{E}\times\hat{b})\Big{)}\Big{]}$
$+\frac{(\omega_{i}-{\rm b}_{i}h)\tau_{R}^{i}}{T(1+\chi_{i})(1+\chi_{i}+\chi_{i}^{2})}\Big{[}(1+\chi_{i})\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{T}\Big{)}$
$-\tau_{R}^{i}\frac{(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{\dot{T}}\Big{)}\pm\chi_{i}(1+\chi_{i})\Big{(}\vec{v_{i}}\cdot(\vec{\nabla}{T}\times\hat{b})\Big{)}$
$\mp\tau_{R}^{i}\frac{\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{(}\vec{v_{i}}\cdot(\vec{\nabla}{\dot{T}}\times\hat{b})\Big{)}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (40)
@@
$j^{l}$ $=\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}|\vec{k}_{i}|}{(2\pi)^{3}}v_{i}^{2}\frac{\tau_{R}^{i}}{(1+\chi_{i}+\chi_{i}^{2})(1+\chi_{i})}\Big{[}-q_{i}\Big{\{}\Big{(}(1+\chi_{i})+\frac{\chi_{i}(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\Big{)}\delta^{lk}E^{k}\pm\Big{(}\chi_{i}(1+\chi_{i})+\frac{\chi_{i}^{2}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{)}$
$\epsilon^{ljk}h^{j}E^{k}\Big{\}}+\frac{(\omega_{i}-{\rm b}_{i}h)}{T}\Big{\{}(1+\chi_{i})\delta^{lk}\frac{\partial{T}}{\partial x^{k}}-\tau_{R}^{i}\frac{(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\delta^{lk}\frac{\partial{\dot{T}}}{\partial x^{k}}\pm\chi_{i}(1+\chi_{i})\epsilon^{ljk}h^{j}\frac{\partial{T}}{\partial x^{k}}\mp\tau_{R}^{i}\frac{\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}$
$\epsilon^{ljk}h^{j}\frac{\partial{\dot{T}}}{\partial x^{k}}\Big{\}}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (41)
@@
$j^{x}=$ $\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}|\vec{k}_{i}|}{(2\pi)^{3}}v_{i}^{2}\frac{\tau_{R}^{i}q_{i}}{(1+\chi_{i}+\chi_{i}^{2})(1+\chi_{i})}\Big{\{}\Big{(}\frac{(1+\chi_{i}^{2})+\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{)}E^{x}\pm\chi\Big{(}\frac{(1+\chi_{i}^{2})(1+\chi_{i})+\chi_{i}(2+\chi_{i})}{(1+\chi_{i}^{2})}\Big{)}(E^{z}-E^{y})\Big{\}}$
$\Big{(}-\frac{\partial f^{0}_{i}}{\partial\omega_{i}}\Big{)}+\sum_{i}\frac{q_{i}g_{i}}{3}\int\frac{d^{3}|\vec{k}_{i}|}{(2\pi)^{3}}v_{i}^{2}\frac{\tau_{R}^{i}(\omega_{i}-{\rm b}_{i}h)}{T(1+\chi_{i}+\chi_{i}^{2})(1+\chi_{i})}\Big{\{}(1+\chi_{i})\frac{\partial T}{\partial x}-\tau_{R}^{i}\frac{(1+\chi_{i}-\chi_{i}^{2})}{(1+\chi_{i}^{2})}\frac{\partial\dot{T}}{\partial x}\pm\chi_{i}(1+\chi_{i})\Big{(}\frac{\partial T}{\partial z}-\frac{\partial T}{\partial y}\Big{)}$
@@
we can write Eq. (IV.1.1) as
$j^{x}$ $=\sum_{i}q_{i}^{2}L_{1i}E_{x}-\sum_{i}q_{i}^{2}L_{2i}E_{y}+\sum_{i}q_{i}^{2}L_{2i}E_{z}-\frac{1}{T}\sum_{i}q_{i}\Big{(}L_{3i}-L_{4i}\frac{d}{dt}\Big{)}\frac{dT}{dx}+\frac{1}{T}\sum_{i}q_{i}\Big{(}L_{5i}-L_{6i}\frac{d}{dt}\Big{)}\frac{dT}{dy}$
$-\frac{1}{T}\sum_{i}q_{i}\Big{(}L_{5i}-L_{6i}\frac{d}{dt}\Big{)}\frac{dT}{dz}~{}.$ (44)
@@
Where, $\sigma_{el}$ = $\sum_{i}q_{i}^{2}L_{1i}$ and $\sigma_{H}$ = $\sum_{i}q_{i}^{2}L_{2i}$ are the Ohmic and the Hall components of electrical conductivity, respectively. Further, we defined the integrals $L_{34}$ = $\sum_{i}q_{i}\Big{(}L_{3i}-L_{4i}\frac{d}{dt}\Big{)}$ and $L_{56}$ = $\sum_{i}q_{i}\Big{(}L_{5i}-L_{6i}\frac{d}{dt}\Big{)}$. To solve the Eqs. (45), we rearrange them in the matrix form of
$\mathbf{\bm{\sigma}~{}E=L~{}X},$ (46)